OpenStudy (carlj0nes):

The cable of a hoist has a cross section of 80 mm^2. The hoist is used to lift a crate weighing 500 kg. The free length of the cable is 3 m. What is the stress in the cable? How much will it extend if it is made of steel (modulus 200 GPa)? How much if it is made of polypropylene, PP (modulus 1.2 GPa)?

OpenStudy (irishboy123):

for first part: \(\sigma = \dfrac{500* g}{80 mm^2}\) sort out units!! for second part: \( \epsilon = \dfrac{\dfrac{500* g}{80 mm^2}}{Y} \) same story.... sort out units!